17) Consider the following reaction. How many moles of oxygen are required to produce 10.0 moes of 17) water? Assume that there is excess C3H7SH present. C3H7SH)+602(g) 3CO2(g)+S02(g)+4 H20(g) A) 15.0 moles O2 B) 6.67 moles O2 C) 10.0 moles O2 D) 40.0 moles O2 E) 250 moles O2 18) How many molecules of HCl are formed when 80.0 g of water reacts according to the following) balanced reaction? Assume excess ICl3. A) 4.46 1024 molecules HCI B) 536 x1024 molecules HCI C)8.04 1024 molecules HC D) 2.68 1024 molecules HCI E) 4.46 1025 molecules HCI 19) Carbonic acid can form water and carbon dioxide upon heating. How much carbon dioxide is 19) formed from 24.8 g of carbonic acid? H2CO3H20 CO2 A) 7.20 g B) 349 g C)248 g D) 17.6g E) 854 g 20) How many grams of calcium chloride are needed to produce 1.50g of potassium chloride? 20) CaCl2(aq) K2CO3(aq) 2 KC(a) CaCco3(aq) A) 0.896 g B) 1.12 g 4.47 g D) 223g 21) If the density of ethanol, C2H5OH, is 0.789 g/mL. How many milliliters of ethanol are needed to21) produce 15.0 g of CO2 according to the following chemical equation? C2H5OH0) 302(g)--2 Co2(g) 3 H200) A) 9.95 ml B) 6.19 mL C)19.9 ml D) 39.8 mL 17-21
Answer #1

17) from equation: 1 mol C3H7SH = 6 mol O2 = 4 mol water

No of mol of O2 required = 10*6/4 = 15 mol

answer: A

18) from equation: 2 mol ICl3 = 3 MOL H2O = 5 mol HCl

no of mol of water = 80/18 = 4.44 mol

no of mol of HCl formed = 4.44*5/3 = 7.4 mol

molecules of HCl = 7.4*6.023*10^23 = 4.46*10^24 molecule

answer: A

19) no of mol of carbonic acid = 24.8/62= 0.4 mol

    no of mol of co2 = 0.4 mol

   amount of co2= 0.4*44= 17.6 g

answer: d

20) no of mol of KCl = 1.5/74.55 = 0.02 mol

    no of mol of cacl2 = 0.02/2 = 0.01 mol

amount of cacl2 = 0.01*111 = 1.11 g

answer: b